OpenStudy (anonymous):

I don't know if I can do this or not, but is there someone good at trig and give give me some study practice problems to do? My teacher gave us a packet to do over the break and I finished it the day we got it, but I would really like soomeone to give some problems to do.

OpenStudy (dtan5457):

Any chance you could get a barron's review book? Got mine from a library a few days ago. It's REALLY good review.

OpenStudy (anonymous):

I have no clue what that is. But I'm googling it right now. I'm taking a Math3 class. I live in NC and they took away Geometry, and Alg. 1 & 2. So we don't have a lot of knowledge on some things. I'm a staight A student in Math but I want some help. I found their website and I'm checking it out right now.

OpenStudy (dtan5457):

It's about $14, tons of fun

OpenStudy (anonymous):

I saw. I have a little bit of money. So I'll talk to my mom to see if she can take me to get it some time.

OpenStudy (dtan5457):

What are you up to on trig?

OpenStudy (anonymous):

We have already done finding angle measures and side lengths. When we go back on the 5th I think, we are going to review law of sine and cosine.

OpenStudy (dtan5457):

Is that past functions, inverses, and complex numbers?

OpenStudy (anonymous):

Umm, hang on let me look I have my binder from last year I can look in. I can't remember off the top of my head.

Directrix (directrix):

Check out this online review Trig Site. @amanda7414 It will check your answers and provide commentary on any you missed. I tested it and it works. http://mathonline.missouri.edu/mucgi-bin/munew.cgi?variable=trig

OpenStudy (dtan5457):

Yes^ I can give you problems if you want, or you can go do them at a website.

OpenStudy (anonymous):

For law of sine it's something like \[\frac{ \sin A }{ a }=\frac{ \sin B }{ b }=\frac{ \sin C }{ c }\] and you're sorta finding angle measures and side lengths using that equation. And for law of cosine \[a ^{2}=b ^{2}+c ^{2}-2bc \cos A\] \[b ^{2}=a ^{2}+c ^{2}-2ac \cos B\] \[c ^{2}=a ^{2}+b ^{2}-2ab \cos C\] and thanks @Directrix

Directrix (directrix):

These self-tests look promising. Answers at the end of each module. http://shawtlr.net/Trigonometry/Self-Tests.html

Directrix (directrix):

Oh, so you want a study guide and then maybe practice problems?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Dear gosh, this is college trig I should be learning this next semester in AFM but this makes no sense to me but I am determined to figure it out.

Directrix (directrix):

Focus on one thing at the time. This link is to "pure" Trig. Now, Algebra permeates Trig and everything else. That's why we study it first. lol Click on the Trig tab. http://mathonline.missouri.edu/mucgi-bin/munew.cgi?variable=trig

OpenStudy (anonymous):

Lol, yeah thanks again I'll get on it

Directrix (directrix):

Look at the attachment It is Trig.

OpenStudy (dtan5457):

@amanda7414 I'll tag you on questions that I can't figure out. Will give you extra practice here and there :)

OpenStudy (anonymous):

ok thanks

Directrix (directrix):

I'm looking for a summary of topics. There are plenty of people here 7 by 24 so post questions that give you fits. And, don't start out assuming you don't know anything, okay. If you don't want 75 questions, set the tab to fewer.

Directrix (directrix):

You are welcome. Whatever else I find, I'll post in this thread.

OpenStudy (anonymous):

lol ok, I'm one to give up until I can figure the problem out

OpenStudy (anonymous):

Ok, so the first problem I haven't learned http://mathonline.missouri.edu/mucgi-bin/munew.cgi I don't understand what MON=pi/4

Directrix (directrix):

We don't give up. Look at this. It's a list of formulas and facts. Do not try to stuff that into your head. http://tutorial.math.lamar.edu/pdf/Trig_Cheat_Sheet.pdf

Directrix (directrix):

The best way to prepare is not to study formula charts but to work problems.

OpenStudy (anonymous):

This is stuff I have not learned but I am determined to learn it and show my teacher I know how to do it, it might take some time but I am going to do it.

Directrix (directrix):

Start with this one: http://mathonline.missouri.edu/mucgi-bin/munew.cgi?variable=trig

Directrix (directrix):

This is my first Missouri question.

OpenStudy (anonymous):

I have something totally different

Directrix (directrix):

That is because of the random generator for the order of the problems.

OpenStudy (anonymous):

oh

Directrix (directrix):

So, that question will turn up with a different number on my test.

OpenStudy (anonymous):

See I never learned pi radians. We have done things in degrees and have never done this I'm taking a basically Alg. 2 but we never learned the unit circle or anything our exam is at the end of January and we haven't even finished learning everything yet.

Directrix (directrix):

If you are in Algebra 2, then you may be doing right triangle trig. I thought you had done a complete trig course. So, I don't know what you have done. You can go through the Missouri test and do the problems over topics you have studied.

OpenStudy (anonymous):

No, I've only done right triangle trig lol, should've been clear on that. That's probably why I'm lost

Directrix (directrix):

Then, those study guides that I posted are not for you at this point in your study.

OpenStudy (anonymous):

Lol, but it gives me a little head start

Directrix (directrix):

I think I would go back through and study the text and practice working problems from the text. And, the re-work the problems the teacher gave you. If you have a problem from Law of Sines or Law of Cosines that you don't understand, then post the problem here. People will help. I always study math by working problems.

OpenStudy (anonymous):

Ok, sounds good and thanks again.